Suppose Acap Corporation will pay a dividend of $ 2.78 per share at the end of this year and $ 2.94 per share next year. You expect​ Acap's stock price to be $ 50.56 in two years. Assume that​ Acap's equity cost of capital is 11.7 %. a. What price would you be willing to pay for a share of Acap stock​ today, if you planned to hold the stock for two​ years? b. Suppose instead you plan to hold the stock for one year. For what price would you expect to be able to sell a share of Acap stock in one​ year? c. Given your answer in ​(b​), what price would you be willing to pay for a share of Acap stock today if you planned to hold the stock for one​ year? How does this compare to your answer in ​(a​)?
Answer #1

a.

Stock Price today = 2.78/(1.117) + (2.94 + 50.56)/(1.117)2

Stock Price today = $45.37

b.

Stock Price in 1 year = (2.94 + 50.56)/(1.117)

Stock Price in 1 year = $47.90

c.

Stock Price = (2.78 + 47.90)/1.117

Stock Price = $45.37